OpenStudy (helpmeeeeeee):

Evaluate 20 + 16 ÷ 2 − 5. 13 18 14 23Evaluate 20 + 16 ÷ 2 − 5. 13 18 14 23

OpenStudy (helpmeeeeeee):

hahaha lol it copied twice

OpenStudy (omw2fyb):

13

OpenStudy (mjazzw7):

The answer is 13. Hope I helped :) PLEASE fan and medal. Thanks

OpenStudy (mjazzw7):

20 + 16 = 36 / 2 = 18 - 5 = 13